# Instruction-Following Pruning for Large Language Models

### v2

@@ Abstract
Instruction-Following Pruning for Large Language Models
Bairu Hou1,2thanks: Work done while interning at Apple. Correspondence to: Bairu Hou $<$bairu@ucsb.edu$>$ and Tao Lei $<$tao_lei2@apple.com$>$. Qibin Chen1 Jianyu Wang1 Guoli Yin1 Chong Wang1Nan Du1Ruoming Pang1Shiyu Chang2Tao Lei1 1Apple AI/ML  2UC Santa Barbara
@@ 1 Introduction
Can LLMs learn to select the most suited parameters based on the task description?
We aim to automatically generate input-specific pruning masks tailored to the tasks described in user prompts. This dynamic, context-aware pruning mechanism enables the language model to perform inference using only the parameters necessary for the task, offering a compelling balance between efficiency and expressivity compared to using a static dense model. Moreover, because the parameters are selected and fixed, our method avoids reloading new parameters during the decoding process. This design choice contrasts with other dynamic methods such as contextual sparsity (Liu et al., 2023; Zhou et al., 2024) and mixture-of-experts Lepikhin et al. (2020); Fedus et al. (2022); Dai et al. (2024), which load different parameters at each decoding step, leading to significant weight loading costs.
Figure 1: Overview of IFPruning. (Left) For each given prompt, the sparsity predictor (much smaller than the LLM) determines which rows and columns of the FFN matrices should be activated. (Middle) The LLM then activates only the selected FFN parameters along with other parameters to perform inference for that specific prompt. (Right) By pruning an 9B LLM to 3B for each input, IFPruning significantly outperforms the dense 3B model and achieves performance levels close to the dense 9B model.
To address the central question, we present Instruction-Following Pruning (IFPruning), a method that integrates a sparse mask predictor with the language model to dynamically generate input-dependent pruning masks, as illustrated in Figure 1. Specifically, we focus on structured pruning of the feed-forward neural network layers, where entire rows or columns are pruned (Xia et al., 2024; Gunter et al., 2024; Sreenivas et al., 2024). The user prompt is first passed into the sparsity predictor, which assigns importance scores to the rows and columns of each feed-forward network layer. These scores are then transformed into differentiable masks using the SoftTopK operator (Ainslie et al., 2023), to achieve a predefined number of sparsity (e.g., reducing a 9B language model to 3B active parameters). The resulting masks are applied to the language model, in which the feed-forward layers are pruned using the masks.
@@
We validate IFPruning through comprehensive experiments across diverse tasks. Specifically, we fine-tune pre-trained language models of varying sizes (6B, 9B, and 12B parameters) using IFPruning and prune them to activate only 3B parameters. In particular, IFPruning consistently outperforms 3B dense models across tasks such as math, coding, tool use, MMLU Hendrycks et al. (2021a) and AlpacaEval (Dubois et al., 2024). For example, when dynamically pruning the 9B model to 3B, our method improves over the 3B dense model by 8% on coding tasks and by 5% on math benchmarks, incurring only marginal performance degradation compared to the unpruned 9B model.
We conduct further analysis to better understand the pruning decisions. Specifically, we observe that instructions requiring similar skills or domain knowledge yield highly homogeneous pruning patterns. Inspired by this analysis, we explore per-task pruning, where a single task prompt generates shared masks for all test instances within the same task. Results show that per-task pruning maintains robust performance while further reducing data loading overhead.
@@ 2 Related Work
Pruning has been extensively studied to compress neural networks and improve their efficiency (Han et al., 2015; Zhu and Gupta, 2017). Previous work has explored different pruning techniques for both unstructured pruning  (Narang et al., 2017; Frankle and Carbin, 2018; Li et al., 2020; Chen et al., 2020) and structured pruning (Wen et al., 2016; Voita et al., 2019; Louizos et al., 2018; Wang et al., 2020). As structured pruning removes entire components in the model such as channels and attention heads, it is more hardware-friendly than unstructured pruning to compress the large models.
Various methods have been proposed for structured pruning of LLMs. LLM-Pruner (Ma et al., 2023) adopt the gradient information to find unimportant components in LLMs and remove them. ShortGPT (Men et al., 2024) proposes to identify and remove those less important layers, where the layer importance is measured by the similarity between inputs and outputs of that layer. In comparison, other optimization-based methods directly learn the parameter masks. For example, Sheared LLaMA  (Xia et al., 2024) use the HardConcrete masking (Louizos et al., 2018; Wang et al., 2020) to generate differentiable masks and optimize the model and masks on pre-training data. Our method also directly optimize the sparsity predictor and the LLM, and we further extend static pruning to input-dependent pruning.
@@
Our approach is also directly motivated by the contextual sparsity of LLMs Liu et al. (2023); Akhauri et al. (2024); Lee et al. (2024). Previous work has identified the existence of input-dependent sub-networks (e.g., attention heads and MLP parameters) within ReLU-based LLMs that can generate approximately the same output as the full model for an input. By predicting such sparsity patterns at each decoding step, we can achieve a favorable balance between accuracy and speedup. But state-of-the-art LLMs (Dubey et al., 2024; Liu et al., 2024a; Yang et al., 2024) design MLP blocks based on more complex non-linear activation functions such as SwiGLU (Shazeer, 2020) that do not inherently induce sparsity (Mirzadeh et al., 2023; Song et al., 2024). Therefore, directly predicting the sparsity patterns can lead to significant performance degradation (Zhou et al., 2024; Dong et al., 2024). In comparison, we co-optimize the sparsity predictor and the LLM with non-ReLU activation functions to achieve better contextual sparsity with minimum performance degradation. Also, most contextual sparsity methods require predicting sparsity and loading different parameters at each decoding step. Our method eliminates this overhead by selecting the parameters based on the input or task description and fixing them for the entire decoding process, avoiding the parameter reloading cost.
@@
Mixture-of-Experts (MoE) have emerged as a popular architecture for scaling LLMs while managing inference costs (Lepikhin et al., 2020; Fedus et al., 2022; Zhou et al., 2022; Dai et al., 2024; Liu et al., 2024b). These models organize every FFN layer into multiple large FFN blocks referred to as experts, and selectively activate a few experts for each input token via a routing mechanism. Compared to MoE, our method selects and fixes the activated parameters given the input task prompt. Although this choice loses the flexibility of using different parameters per token, it significantly reduces weight loading costs for decoding. In addition, our method performs much fine-grained selection of parameters by activating or pruning each FFN dimension independently, which enhances model expressivity.
@@ 3.1 Overview of Structured Pruning
Denote the hidden dimension of the LLM as $d$, the intermediate dimension of the FFN blocks as $d_{\mathrm{ffn}}$, the input length as $n$, and $X\in\mathbb{R}^{n\times d}$ as the input of a transformer FFN block $F_{\mathrm{ffn}}(\cdot)$. The goal of structured pruning is to reduce FFN intermediate dimension from $d_{\mathrm{ffn}}$ to $t_{\mathrm{ffn}}$. Without loss of generality, consider a standard FFN block defined as
@@ 3.2 Architecture
The sparsity predictor consists of two modules: ❶ a much smaller LLM backbone to extract the features of user prompts and ❷ a mask prediction head. Specifically, the LLM backbone takes the prompt $\bm{x}=(x_{1},\dots,x_{n})$ as input with length $n$, and we use the hidden states of the last token $x_{n}$ in the last layer to represent the prompt. The mask prediction head is a two-layer MLP, which predicts the masks given the prompt presentations. The output of the FFN mask prediction head is the masking score $\bm{z}\in\mathbb{R}^{L\times d_{\mathrm{ffn}}}$, where $L$ is the number of layers of the LLM.
Given the predicted masking score $\bm{z}$, a mask generation operator will be applied to $\bm{z}$ to convert it to the mask $\bm{m}\in[0,1]^{L\times d_{\mathrm{ffn}}}$, which contains $t_{\mathrm{ffn}}$ nonzero elements. In this paper, we use the SoftTopK (Lei et al., 2023; Ainslie et al., 2023) algorithm to generate a differentiable $\bm{m}$, but we also acknowledge that other algorithms such as the HardConcrete masking Louizos et al. (2018); Wang et al. (2020) are also applicable. Particularly, given the FFN masking score $\bm{z}$, SoftTopK converts it to masks $\bm{m}$ via:
@@
Here $\bm{z}^{(i)}$, $\bm{\lambda}^{(i)}$ and $\bm{m}^{(i)}$ represent the $i$-th row of each matrix, $g(\cdot):\mathbb{R}^{d_{\mathrm{ffn}}}\rightarrow[0,1]^{d_{\mathrm{ffn}}}$ is a normalization function, and $\mathrm{Top}(\cdot,t_{\mathrm{ffn}})\in\{0,1\}^{d_{\mathrm{ffn}}}$ is an indicator function that returns a binary mask indicating the top-k values in $\lambda$. The normalization function $g(\cdot)$ ensures that $\lambda$ satisfies the sparsity constraint, i.e., $\sum_{k}\lambda^{(i)}_{k}=t_{\mathrm{ffn}}$, where $t_{\mathrm{ffn}}$ is the target size of the FFN layers. More details of SoftTopK can be found in the previous work Lei et al. (2023); Ainslie et al. (2023).
@@
During training, the LLM takes the masks $\bm{m}$ as an input and prune its FFN blocks. We use standard next token prediction loss computed over tokens within a training batch, and we co-optimize the LLM and the sparsity predictor.
@@ 3.3 Model Training
where $\bm{\theta}$ refers to the parameters of the LLM, $\bm{m}^{(k)}$ is the predicted mask based on chunk $\bm{x}^{(k)}$, $f(\bm{x}_{<i};\bm{\theta},\bm{m}^{(k)})$ is the next token prediction distribution from the LLM with $\bm{m}^{(k)}$ applied, and $\ell(\cdot)$ is the Cross-Entropy loss. Since the chunks are consecutive, the sparsity predictor can learn to utilize the contextual information of each chunk to predict which parameters of the LLM are best suited for the next token prediction. Because both the sparsity predictor and LLM are co-optimized in this stage, it provides a good initialization for the fine-tuning stage.
@@
Starting from the models after the first stage, we train the sparsity predictor and the LLM on a supervised fine-tuning dataset that contains several million examples. Our SFT data contains a diverse set of prompts to predict the sub-networks. Some prompts specify the task and an input, while others include few-shot examples along with the input. Lastly, many examples only contain a task description, like “translate English text to French”.
@@ 4.1 Experiment Setup
We compare our method with the following models: ❶ Dense-3B. We compare with a 3B dense LLM that is trained using twice as many pretraining tokens compared to our models. ❷ Pruning+Distill. We also compare our method with static pruning approaches. Similar to recent work such as Sheared LLaMA (Xia et al., 2024), LLAMA 3.2 (Meta AI, 2024), and Minitron(Sreenivas et al., 2024), we include a baseline where a 3B dense LLM is pruned and distilled from a larger pre-trained LLM. we first prune the larger LLM into a dense 3B model by learning masks on the FFN layers similar to Sheared LLaMA (Xia et al., 2024). After pruning, the model undergoes further continuous pre-training through knowledge distillation (Hinton, 2015), using the larger LLM as the teacher model. Therefore, this approach serves as a stronger baseline compared to models using pruning alone. ❸ Dense-9B. We also include a 9B dense model without pruning as a reference for upper-bound performance.
@@
Table 1: Performance comparison between IFPruning and other models. The best results are highlighted in bold and the second-best results are underlined. Dense-9B is included for reference.
@@
  • Instruction-following. We include IFEval (Zhou et al., 2023), AlpacaEval 2.0 (Dubois et al., 2024), and Arena Hard (Li et al., 2024) for evaluation. We report the prompt-level (IFEval-P) and instruction-level (IFEval-I) accuracy for IFEval, the length-controlled win rate for AlpacaEval 2.0, and win rate for Arena Hard.
  • Coding. We evaluate the pass@1 on HumanEval-python (Chen et al., 2021), mbpp (Austin et al., 2021), and MultiPL-E (Cassano et al., 2022). For MultiPL-E benchmark, we use the Swift subset for evaluation.
@@
We mainly use LM-Evaluation-Harness (Gao et al., 2021) to evaluate the tasks, with the exception of instruction-following and tool-use tasks, which are based on the official implementations.
Figure 2: Scaling behavior of dense models and IFPruning. IFPruning activates 3B model for each input. The x-axis represents the total number of LLM parameters for dense models and IFPruning, while the y-axis indicates performance scores on evaluation benchmarks.

### v3

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Can LLMs learn to select the most suited parameters
based on the task description?
We aim to automatically generate input-specific pruning masks tailored to the tasks described in user prompts. This dynamic, context-aware pruning mechanism enables the language model to perform inference using only the parameters necessary for the task, offering a compelling balance between efficiency and expressivity compared to using a static dense model. Moreover, because the parameters are selected and fixed, our method avoids reloading new parameters during the decoding process. This design choice contrasts with other dynamic methods such as contextual sparsity (Liu et al., 2023; Zhou et al., 2024) and mixture-of-experts (Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024), which load different parameters at each decoding step, leading to significant weight loading costs. Our design is particularly suited for on-device models (e.g. smartphones and laptops), where the inference typically samples a few responses given the same user query (or the same task). In this case, the same activated parameters are selected and cached as a dense model, therefore achieving the same speedup as the static pruning.
Figure 1: Overview of IFPruning. (Left) For each given prompt, the sparsity predictor (much smaller than the LLM) determines which rows and columns of the FFN matrices should be activated. (Middle) The LLM is then pruned accordingly and uses the selected parameters to perform inference for that specific prompt. (Right) By pruning a 9B LLM to 3B for each input, IFPruning significantly outperforms the dense 3B model and achieves performance levels close to the dense 9B model. It also achieves nearly the same inference latency as the dense 3B model, as measured by time-to-first-token (TTFT).
To address the central question, we present Instruction-Following Pruning (IFPruning), a method that integrates a sparsity predictor with the language model to dynamically generate input-dependent pruning masks, as illustrated in Figure 1. Specifically, we focus on structured pruning of the feed-forward neural network layers, where entire rows or columns of the weight matrices are pruned (Xia et al., 2024; Gunter et al., 2024; Sreenivas et al., 2024). The user prompt is first passed into the sparsity predictor, which assigns importance scores to the rows and columns of each feed-forward network layer. These scores are then transformed into differentiable masks using the SoftTopK operator (Ainslie et al., 2023), to achieve a predefined number of sparsity (e.g., reducing a 9B language model to 3B active parameters). The resulting masks are applied to the language model, in which the feed-forward layers are pruned using the masks.
@@
We validate IFPruning through comprehensive experiments across diverse tasks in Section 4. Specifically, we fine-tune pre-trained language models of varying sizes (6B, 9B, and 12B parameters) using IFPruning and prune them to activate only 3B parameters. In particular, IFPruning consistently outperforms 3B dense models across tasks such as math, coding, tool use, MMLU (Hendrycks et al., 2021a) and AlpacaEval (Dubois et al., 2024). For example, when dynamically pruning the 9B model to 3B, our method improves over the 3B dense model by 8% on coding tasks and by 5% on math benchmarks, incurring only marginal performance degradation compared to the unpruned 9B model.
We conduct further analysis to better understand the pruning decisions. Specifically, we observe that instructions requiring similar skills or domain knowledge yield highly homogeneous pruning patterns. Inspired by this analysis, we explore per-task pruning in Section 4.3, where a single task prompt generates shared masks for all test instances within the same task. Results show that per-task pruning maintains robust performance while further reducing data loading overhead.
We also show that IFPruning can significantly improve the LLM inference efficiency in Section 4.4. Compared to the full model, IFPruning reduces the time-to-first token by up to 57% and the generate time by up to 41%. In addition, the overhead introduced by dynamic pruning and parameter caching is negligible, adding less than 0.1 seconds per example and accounting for only 1–2% of the total generation time.
@@ 2 Related Work
In this section, we provide an overview of prior research that closely relates to and partially motivates our work, including model pruning, contextual sparsity, and mixture-of-experts.
@@
Pruning has been extensively studied to compress neural networks and improve their efficiency (Han et al., 2015; Zhu & Gupta, 2017). Previous work has explored different pruning techniques for both unstructured pruning  (Narang et al., 2017; Frankle & Carbin, 2018; Li et al., 2020; Chen et al., 2020) and structured pruning (Wen et al., 2016; Voita et al., 2019; Louizos et al., 2018; Wang et al., 2020). As structured pruning removes entire components in the model such as channels, attention heads, and feed-forward neural network intermediate dimensions, it is more hardware-friendly than unstructured pruning to compress the large models.
Various methods have been proposed for structured pruning of LLMs (Yang et al., 2024b; Kim et al., 2024; Kurtić et al., 2024; Dery et al., 2024). LLM-Pruner (Ma et al., 2023) adopt the gradient information to find unimportant components in LLMs and remove them. SliceGPT transforms each weight matrix in transformer blocks into a smaller one by applying orthogonal transformations to reduce the embedding dimensions of weight matrices. ShortGPT (Men et al., 2024) proposes to identify and remove those less important layers, where the layer importance is measured by the similarity between inputs and outputs of that layer. In comparison, other optimization-based methods directly learn the parameter masks. For example, Sheared LLaMA  (Xia et al., 2024) use the HardConcrete masking (Louizos et al., 2018; Wang et al., 2020) to generate differentiable masks and optimize the model and masks on pre-training data. Our method also directly optimize the sparsity predictor and the LLM, and we further extend static pruning to input-dependent pruning.
@@
Our approach is also directly motivated by the contextual sparsity of LLMs (Liu et al., 2023; Akhauri et al., 2024; Lee et al., 2024). Previous work has identified the existence of input-dependent sub-networks (e.g., attention heads and MLP parameters) within ReLU-based LLMs that can generate approximately the same output as the full model for an input. By predicting such sparsity patterns at each decoding step, we can achieve a favorable balance between accuracy and speedup. But state-of-the-art LLMs (Dubey et al., 2024; Liu et al., 2024a; Yang et al., 2024a) design MLP blocks based on more complex non-linear activation functions such as SwiGLU (Shazeer, 2020), SiLU (Elfwing et al., 2018; Ramachandran et al., 2017) and GELU (Hendrycks & Gimpel, 2016) that do not inherently induce sparsity (Mirzadeh et al., 2023; Song et al., 2024). Therefore, directly predicting the sparsity patterns can lead to significant performance degradation (Zhou et al., 2024; Dong et al., 2024). In comparison, we co-optimize the sparsity predictor and the LLM with non-ReLU activation functions to achieve better contextual sparsity with minimum performance degradation. Also, most contextual sparsity methods require predicting sparsity and loading different parameters at each decoding step. Our method eliminates this overhead by selecting the parameters based on the input or task description before decoding starts. The selected parameters are fixed for the entire decoding process, avoiding the parameter reloading cost.
@@
Mixture-of-Experts (MoE) have emerged as a popular architecture for scaling LLMs while managing inference costs (Lepikhin et al., 2020; Du et al., 2022; Fedus et al., 2022; Zhou et al., 2022; Dai et al., 2024; Liu et al., 2024b). These models organize every FFN layer into multiple large FFN blocks referred to as experts, and selectively activate a few experts for each input token via a routing mechanism (Lepikhin et al., 2020; Zoph et al., 2022; Sun et al., 2024). Our method share the same spirit as MoE by dynamically activating a subset of parameters. However, our method selects the activated parameters given the input prompt, and reuses the same activated parameters during decoding. Although this choice loses the flexibility of using different parameters per token, it significantly reduces weight loading costs for decoding. In this regard, our model is a sparse model designed for on-device scenarios where both memory and computational resources are constrained. Another difference is that our method performs more fine-grained selection of parameters by activating or pruning each FFN dimension independently, which enhances model expressivity.
@@ 3.1 Overview of Structured Pruning
Denote the hidden dimension of the LLM as $d$, the intermediate dimension of the FFN blocks as $d_{\mathrm{ffn}}$, the input length as $n$, and $X\in\mathbb{R}^{n\times d}$ as the input of a transformer FFN block $F_{\mathrm{ffn}}(\cdot)$. The goal of our structured pruning method is to reduce the FFN intermediate dimension from $d_{\mathrm{ffn}}$ to $t_{\mathrm{ffn}}$. Without loss of generality, consider a standard FFN block defined as
@@ 3.2 Architecture
The sparsity predictor consists of two modules: ❶ a much smaller LLM backbone to extract the features of user prompts and ❷ a mask prediction head. Specifically, the LLM backbone takes the prompt $\bm{x}=(x_{1},\dots,x_{n})$ as input with length $n$, and we use the hidden states of the last token $x_{n}$ in the last layer to represent the prompt. The mask prediction head is a two-layer MLP, which predicts the masks given the prompt presentations. The output of the FFN mask prediction head is the masking score $\bm{z}\in\mathbb{R}^{L\times d_{\mathrm{ffn}}}$, where $L$ is the number of layers of the LLM. We include more details about the architecture of the sparsity predictor in Appendix A.1.
Given the predicted masking score $\bm{z}$, a mask generation operator will be applied to $\bm{z}$ to convert it to the mask $\bm{m}\in[0,1]^{L\times d_{\mathrm{ffn}}}$, which contains $t_{\mathrm{ffn}}$ nonzero elements. In this paper, we use the SoftTopK (Lei et al., 2023; Ainslie et al., 2023) algorithm to generate a differentiable $\bm{m}$, but we also acknowledge that other algorithms such as the HardConcrete masking (Louizos et al., 2018; Wang et al., 2020) are also applicable. Particularly, given the FFN masking score $\bm{z}$, SoftTopK converts it to masks $\bm{m}$ via:
@@
Here $\bm{z}^{(i)}$, $\bm{\lambda}^{(i)}$ and $\bm{m}^{(i)}$ represent the $i$-th row of each matrix, $g(\cdot):\mathbb{R}^{d_{\mathrm{ffn}}}\rightarrow[0,1]^{d_{\mathrm{ffn}}}$ is a normalization function, and $\mathrm{Top}(\cdot,t_{\mathrm{ffn}})\in\{0,1\}^{d_{\mathrm{ffn}}}$ is an indicator function that returns a binary mask indicating the top-k values in $\lambda$. The normalization function $g(\cdot)$ ensures that $\lambda$ satisfies the sparsity constraint, i.e., $\sum_{k}\lambda^{(i)}_{k}=t_{\mathrm{ffn}}$, where $t_{\mathrm{ffn}}$ is the target size of the FFN layers. More details of SoftTopK can be found in the previous work (Lei et al., 2023; Ainslie et al., 2023).
@@
During training, the LLM takes the masks $\bm{m}$ as an additional input and prune its FFN blocks. We use standard next token prediction loss computed over tokens within a training batch, and we co-optimize the LLM and the sparsity predictor.
@@ 3.3 Model Training
where $\bm{\theta}$ refers to the parameters of the LLM, $\bm{m}^{(k)}$ is the predicted mask based on chunk $\bm{x}^{(k)}$, $f(\bm{x}_{<i};\bm{\theta},\bm{m}^{(k)})$ is the next token prediction distribution from the LLM with $\bm{m}^{(k)}$ applied, and $\ell(\cdot)$ is the Cross-Entropy loss. Since the chunks are consecutive, the sparsity predictor can learn to utilize the contextual information of each chunk to predict which parameters of the LLM are best suited for the next token prediction in the next text chunk. Because both the sparsity predictor and LLM are co-optimized in this stage, it provides a good initialization for the fine-tuning stage.
@@
Starting from the models after the first stage, we train the sparsity predictor and the LLM on a supervised fine-tuning dataset that contains several million examples. During training, the input prompts will be fed into the sparsity predictor which predicts the masks for each input. The LLM is then masked and optimized to predict the target outputs conditioned on the input prompts. Our SFT data contains a diverse set of prompts to predict the sub-networks. Some prompts specify the task with a task description and an input, while others include few-shot examples along with the input. Lastly, many examples only contain a task description, like “write a comprehensive blog post about the top 10 most eco-friendly cities in the world.”.
@@ 4.1 Experiment Setup
We compare our method with the following models: ❶ Dense-3B. We compare with a 3B dense LLM that is trained using twice as many pretraining tokens compared to our models. ❷ Pruning+Distill. We also compare our method with static pruning approaches. Similar to recent work such as Sheared LLaMA (Xia et al., 2024), LLAMA 3.2 (Meta AI, 2024), and Minitron(Sreenivas et al., 2024), we include a baseline where a 3B dense LLM is pruned and distilled from a larger pre-trained LLM. we first prune the larger LLM into a dense 3B model by learning masks on the FFN layers similar to Sheared LLaMA (Xia et al., 2024). After pruning, the model undergoes further continuous pre-training through knowledge distillation (Hinton, 2015), using a larger dense model with 12B parameters as the teacher model. Therefore, this approach serves as a stronger baseline compared to models using pruning alone. More details of the knowledge distillation can be found in Appendix A.2. ❸ Dense-9B. We also include a 9B dense model without pruning as a reference for upper-bound performance.
Table 1: Performance comparison between IFPruning and other dense LLMs. Pruning+Distill 3B is first pruned from larger model into 3B parameters and undergoes continuous pre-training with knowledge distillation using a 12B LLM as the teacher model. Three versions of IFPruning are included which activate 3B parameters of LLMs with 6B, 9B, and 12B parameters. The best results are highlighted in bold and the second-best results are underlined. Dense-9B is included for reference.
@@
  • Instruction-following. We include IFEval (Zhou et al., 2023), AlpacaEval 2.0 (Dubois et al., 2024), and Arena-Hard-Auto (Li et al., 2024) for evaluation. We report the prompt-level and instruction-level accuracy on IFEval, the length-controlled win rate on AlpacaEval 2.0, and win rate on Arena-Hard-Auto.
  • Coding tasks. We evaluate the pass@1 performance on HumanEval-python (Chen et al., 2021), mbpp (Austin et al., 2021), and MultiPL-E (Cassano et al., 2022). For MultiPL-E benchmark, we use the Swift subset for evaluation.
@@
We mainly use LM-Evaluation-Harness (Gao et al., 2021) to evaluate the tasks, with the exception of instruction-following and tool-use tasks, which are based on their official implementations.
Figure 2: Scaling behavior of dense models and IFPruning. IFPruning activates 3B parameters for each input. The x-axis represents the total number of LLM parameters for dense models and IFPruning, while the y-axis indicates the performance scores.